# Fast Crystal Tensor Property Prediction: A General O(3)-Equivariant Framework Based on Polar Decomposition

### v2

@@ A.3 Experimental details and additional results
Hyperparameter settings of GoeCTP. When constructing the crystal graph, we used the 16th nearest atom to determine the cutoff radius. For edge embeddings, we used an RBF kernel with $c=0.75$ and values ranging from --4 to 0, which maps $-c/||\mathbf{e}_{ij}||_{2}$ to a 512-dimensional vector. In the dielectric tensor prediction task, the 512-dimensional vector is mapped to a 128-dimensional vector through a non-linear layer, while in the elastic tensor prediction, it is mapped to a 256-dimensional vector. For the prediction module, eComFormer, in the dielectric tensor prediction task, we used 4 node-wise transformer layers and 1 node-wise equivariant updating layer, whereas in the elastic tensor prediction, we used 2 node-wise transformer layers and 1 node-wise equivariant updating layer. For both dielectric and elastic tensor tasks, the learning rate was set to 0.001, with 200 epochs and a batch size of 64. In the Reverse R&R module, for the dielectric tensor prediction task, the 128-dimensional vector features output by the prediction module are mapped to a 6-dimensional vector through two non-linear layers, then reconstructed into a $3\times 3$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ obtained from the R&R module to achieve $O(3)$ equivariant output. For the elastic tensor prediction task, the 256-dimensional vector features output by the prediction module are mapped to a 36-dimensional vector through two non-linear layers, then reconstructed into a $6\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
1st 2nd 3rd 4th 5th mean
MEGNET
Fnorm $\downarrow$ 3.62 3.70 3.66 3.74 3.81 3.71
EwT 25% $\uparrow$ 76.0% 77.1% 75.6% 73.7% 76.4% 75.8 %
EwT 10% $\uparrow$ 41.2% 39.3% 38.2% 39.7% 36.3% 38.9%
EwT 5% $\uparrow$ 17.6% 16.8% 18.5% 19.1% 17.8% 18.0%
ETGNN
Fnorm $\downarrow$ 3.42 3.48 3.33 3.35 3.41 3.40
EwT 25% $\uparrow$ 82.1% 81.7% 83.4% 82.6% 83.2% 82.6 %
EwT 10% $\uparrow$ 46.9% 47.8% 50.7% 49.4% 50.5% 49.1%
EwT 5% $\uparrow$ 22.3% 26.5% 27.1% 25.9% 24.6% 25.3%
GMTNet
Fnorm $\downarrow$ 2.98 3.64 3.23 3.16 3.39 3.28
EwT 25% $\uparrow$ 83.4% 81.3% 83.8% 84.5% 83.7% 83.3 %
EwT 10% $\uparrow$ 54.9% 56.2% 55.8% 56.2% 57.1% 56.0%
EwT 5% $\uparrow$ 31.2% 30.5% 28.8% 29.3% 32.7% 30.5%
GoeCTP (Ours)
Fnorm $\downarrow$ 3.33 3.29 3.36 3.25 2.92 3.23
EwT 25% $\uparrow$ 82.5% 83.4% 82.1% 83.9% 84.2% 83.2 %
EwT 10% $\uparrow$ 57.5% 53.9% 56.4% 56.7% 59.7% 56.8%
EwT 5% $\uparrow$ 36.9% 31.6% 37.1% 34.2% 37.8% 35.5%
Table 10: Comparison of 5 repeated experiments on dielectric dataset.
@@
Hyperparameter settings of GMTNet, ETGNN, and MEGNET. Following Yan et al. (2024b), we trained GMTNet, ETGNN, and MEGNET for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks. The same polynomial learning rate decay scheduler is used in all experiments.
@@
Additional results. Given the relatively small size of the dielectric tensor dataset, along with other influencing factors, we observed some randomness in the results during repeated training runs, introducing noise into the outcome comparisons. To mitigate this effect and ensure the reliability of the results, we conducted five independent trials for each method, using the average of the performance metrics across these five trials as the final reported value. The detailed results from these repeated experiments are presented in Table 10.
We assessed the performance of GoeCTP when combined with iComFormer (Yan et al., 2024a) and CrystalFormer (Taniai et al., 2024) (denoted as GoeCTP (iCom.) for the combination) on both the dielectric tensor and elastic tensor datasets. As presented in Table 11, on the dielectric tensor dataset, although the prediction accuracy of GoeCTP (iCom.) and GoeCTP (Crys.) did not exceed that of GoeCTP (eCom.), GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.). In contrast, as shown in Table 12, on the elastic tensor dataset, GoeCTP (iCom.) and GoeCTP (Crys.) achieved superior prediction quality relative to GoeCTP (eCom.), albeit with lower efficiency.

### v3

@@ A.10 Experimental details
Hyperparameter settings of GoeCTP. When constructing the crystal graph, we used the 16th nearest atom to determine the cutoff radius. For edge embeddings, we used an RBF kernel with $c=0.75$ and values ranging from --4 to 0, which maps $-c/||\mathbf{e}_{ij}||_{2}$ to a 512-dimensional vector. In the dielectric and piezoelectric tensor prediction task, the 512-dimensional vector is mapped to a 128-dimensional vector through a non-linear layer, while in the elastic tensor prediction, it is mapped to a 256-dimensional vector. For the prediction module, eComFormer, in the dielectric and piezoelectric tensor prediction task, we used 4 node-wise transformer layers and 1 node-wise equivariant updating layer, whereas in the elastic tensor prediction, we used 2 node-wise transformer layers and 1 node-wise equivariant updating layer. For both dielectric, piezoelectric, and elastic tensor tasks, the learning rate was set to 0.001, with 200 epochs and a batch size of 64. In the Reverse R&R module, for the dielectric tensor prediction task, the 128-dimensional vector features output by the prediction module are mapped to a 6-dimensional vector through two non-linear layers, then reconstructed into a $3\times 3$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ obtained from the R&R module to achieve $O(3)$ equivariant output. For the piezoelectric prediction task, the 18-dimensional vector features output by the prediction module are reconstructed into a $3\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output. For the elastic tensor prediction task, the 256-dimensional vector features output by the prediction module are mapped to a 36-dimensional vector through two non-linear layers, then reconstructed into a $6\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
@@
Hyperparameter settings of GMTNet, ETGNN, and MEGNET. Following Yan et al. [51], we trained GMTNet, ETGNN, and MEGNET for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks. The same polynomial learning rate decay scheduler is used in all experiments.
A.11 Additional results.
We assessed the performance of GoeCTP when combined with iComFormer [50] and CrystalFormer [42] (denoted as GoeCTP (iCom.) for the combination) on both the dielectric, piezoelectric, and elastic tensor datasets. As presented in Table 12, on the dielectric tensor dataset, although the prediction accuracy of GoeCTP (iCom.) and GoeCTP (Crys.) did not exceed that of GoeCTP (eCom.), GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.). As shown in Table 13, similar to the dielectric tensor dataset, GoeCTP (eCom.) achieved the best prediction accuracy on the piezoelectric dataset, but GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.). In contrast, as shown in Table 14, on the elastic tensor dataset, GoeCTP (iCom.) and GoeCTP (Crys.) achieved superior prediction quality relative to GoeCTP (eCom.), albeit with lower efficiency.
@@
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 0.504 0.515 0.450 0.431 0.448 0.457
EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9% 45.5% 42.7%
EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9% 44.9% 42.1%
EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7% 38.3% 41.5%
Total Time (s) $\downarrow$ 943 1220 5771 938 909 2787
Time/batch (s) $\downarrow$ 0.074 0.095 0.45 0.073 0.071 0.218
Table 13: Additional comparison of performance metrics on piezoelectric dataset.
@@
A.12 The limitations of our method and its broader impact
Limitations. (1)As discussed in the main text, GoeCTP is a plug-and-play $O(3)$-equivariant framework designed to enhance the backbone network’s ability to achieve equivariant predictions. Consequently, the performance of our method is inherently dependent on the capabilities of the backbone network. If the backbone lacks sufficient predictive power, the combined framework may not fully surpass the state-of-the-art tensor prediction networks in all evaluation metrics. For instance, as shown in Table 12, when our framework is integrated with CrystalFormer [42], although the computational speed significantly exceeds that of GMTNet, the Fnorm, EwT 25%, and EwT 10% values are lower than those achieved by GMTNet.
(2)As demonstrated in Higham [12] and Proposition 3.5, polar decomposition applied to a $3\times 3$ invertible matrix $\mathbf{L}$ produces a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$. For 3D crystal structures, the lattice matrix $\mathbf{L}$ is always full-rank (i.e. invertible), therefore, ensuring the applicability of our method to 3D crystal systems. However, for certain special cases, such as 2D crystals with single-layer structures [32, 39], the rank of the lattice matrix $\mathbf{L}$ may be less than 3. In these scenarios, directly applying polar decomposition may not yield a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$ and a unique canonical form $\mathbf{H}$. This limitation could cause our method to fail, as crystals with different spatial orientations cannot be consistently adjusted to a unique canonical form, thereby preventing the method from achieving the desired $O(3)$-equivariance.
Broader impact. This work can be applied to screen existing crystal material databases for potential material candidates, facilitating the discovery of new materials with desirable tensor properties, such as high-dielectric materials. Therefore, our work has the potential to make a significant impact on materials science.
A.13 How to further utilize the tensor properties symmetry
In this section, we will use the dielectric tensor as an example to briefly discuss how to further utilize the symmetry of tensor properties.
Utilizing symmetry constraints for zero elements. To demonstrate the role of symmetry constraints, we present an example of the GoeCTP prediction results in Table 15. Within a reasonable margin of error, the predictions align well with the expected constraints.
Label Prediction Cubic dielectric tensor
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.252&0.016&0.008\\ 0.016&2.230&0.007\\ 0.008&0.007&2.262\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Table 15: An example of GoeCTP prediction results
Crystal system Cubic Tetragonal Hexagonal-Trigonal Orthorhombic Monoclinic
Success rate 88.3% 86.6% 84.5% 84.5% 75.7%
Table 16: The GoeCTP results of predicting symmetry-constrained zero-valued dielectric tensor elements.
For a dielectric tensor, we evaluate the prediction of zero elements by defining a threshold as 1% of the average value of non-zero elements in the labels. Predictions are deemed successful if they meet this threshold. The results are as shown in Table 16. It can be observed that our method predicts most zero elements accurately, though not flawlessly. Since our advantage lies in transferring equivariance through an external framework, there are no restrictions on the model itself. Therefore, to achieve a higher success rate for zero elements, we added a ReLU activation function to the output layer of the network to improve the success rate (this applies only to cases where tensor elements are greater than or equal to zero). The results after retraining GoeCTP are as shown in Table 17, Table 18, and Table 19. Our method predicts zero elements more accurately while maintaining overall performance.
Label Prediction Cubic dielectric tensor
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.237&0.000&0.000\\ 0.000&2.283&0.000\\ 0.000&0.000&2.228\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Table 17: An example of GoeCTP (ReLU) prediction results
Crystal system Cubic Tetragonal Hexagonal-Trigonal Orthorhombic Monoclinic
Success rate 100% 100% 87.2% 100% 100%
Table 18: The GoeCTP (ReLU) results of predicting symmetry-constrained zero-valued dielectric tensor elements.
GoeCTP GoeCTP (ReLU)
Fnorm $\downarrow$ 3.23 3.26
EwT 25% $\uparrow$ 83.2% 82.6%
EwT 10% $\uparrow$ 56.8% 58.4%
EwT 5% $\uparrow$ 35.5% 36.3%
Table 19: Performance comparisons between GoeCTP and GoeCTP (ReLU) on the dielectric dataset.
This simple modification enables GoeCTP to more accurately predict the zero elements in dielectric tensors caused by the space group constraints.
Utilizing symmetry constraints for non-zero elements. The symmetry constraints for zero elements serve as a foundational example. When the space group of the input crystal is known, prior knowledge (refer to Appendix A.9) can be employed to ensure that network outputs fully adhere to tensor property constraints. There are two specific methods to achieve this.
In the first method, only the independent components of the tensor properties are predicted. For example, for a cubic crystal, we predict only one independent component and then use the mask from Table 8 to reconstruct it into a $3\times 3$ tensor). In Table 15, GoeCTP was modified to directly predict independent components. The new results are shown in Table 20, which fully satisfy the symmetry constraints. As shown in Table 21, the overall prediction performance indicates that directly predicting the independent components improves the model’s Fnorm performance on the dielectric dataset. Since this method directly predicts independent components, it is more suitable for cases like the dielectric tensor, where the independent components exhibit relatively simple structures. However, for more complex datasets, such as those involving piezoelectric and elastic tensors, the model’s performance using this method may degrade. In such scenarios, the second method provides a more viable alternative.
Label Prediction Cubic dielectric tensor
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.357&0.000&0.000\\ 0.000&2.357&0.000\\ 0.000&0.000&2.357\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Table 20: An example of GoeCTP (predicting the independent components) prediction results
GoeCTP GoeCTP (independent)
Fnorm $\downarrow$ 3.23 3.11
EwT 25% $\uparrow$ 83.2% 80.5%
EwT 10% $\uparrow$ 56.8% 57.3%
EwT 5% $\uparrow$ 35.5% 35.5%
Table 21: Performance comparisons between GoeCTP and GoeCTP (predicting the independent components) on the dielectric dataset.
In the second method, once the network training is completed, the network output is weighted using the standard tensor property format, referred to as a mask, as detailed in Table 8. The application of this method to the piezoelectric and elastic datasets is demonstrated in Tables 22 and 23. The results indicate that incorporating symmetry-based weighting effectively improves the performance of GoeCTP.
GoeCTP GoeCTP (weighted)
Fnorm $\downarrow$ 0.431 0.389
EwT 25% $\uparrow$ 46.9% 48.5%
EwT 10% $\uparrow$ 42.9% 46.7%
EwT 5% $\uparrow$ 39.7% 46.1%
Table 22: Performance comparisons between GoeCTP and GoeCTP (weighted by the standard tensor property form) on the piezoelectric dataset.
GoeCTP GoeCTP (weighted)
Fnorm $\downarrow$ 107.11 106.24
EwT 25% $\uparrow$ 42.5% 46.3%
EwT 10% $\uparrow$ 15.3% 16.5%
EwT 5% $\uparrow$ 7.2% 7.4%
Table 23: Performance comparisons between GoeCTP and GoeCTP (weighted by the standard tensor property form) on the elastic dataset.
A.14 Analysis of discontinuity in canonicalization
Previous studies [6] have shown that canonicalization can often introduce discontinuities into the model. To assess the impact of such discontinuities on our method, we analyze our proposed method from both theoretical and empirical perspectives.
From theoretical perspective: The following proposition, presented in Appendix B of Dym et al. [6], characterizes the existence of a continuous canonicalization under orthogonal group actions:
Proposition A.3.
Consider the action of the orthogonal group $O(d)$ on $\mathbf{X}\in\mathbb{R}^{n\times d}$. There exists a continuous canonicalization for $O(d)$ when $n\leq d$. This continuous canonicalization can be defined as $y_{\mathbf{X}}=\left(\mathbf{X}^{\top}\mathbf{X}\right)^{1/2}$, where $\mathbf{X}^{\top}\mathbf{X}$ forms a positive semi-definite matrix and the square root is the standard square root of a positive semi-definite matrix, i.e. $\left(\mathbf{A}\right)^{1/2}$ is the unique positive semidefinite matrix $\mathbf{B}$ such that $\mathbf{B}^{2}=\mathbf{A}$.
In fact, the polar decomposition adopted in this work can be regarded as a case of such continuous canonicalization.
Proof.
According to Proposition A.3, for a matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$, its continuous canonicalization norm is given by $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}$.
In this work, the polar decomposition is applied to the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H},\mathbf{L}\in\mathbb{R}^{3\times 3}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{H}$ is the positive semi-definite Gram matrix (which serves as the canonicalization norm). Based on the properties of positive semi-definite matrices, our method is actually equivalent to the canonicalization approach described in Proposition A.3, i.e. $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}=\left((\mathbf{Q}\mathbf{H})^{\top}\mathbf{Q}\mathbf{H}\right)^{1/2}=\left(\mathbf{H}^{\top}\mathbf{H}\right)^{1/2}=\mathbf{H}$. ∎
Therefore, our canonicalization via polar decomposition is consistent with the continuous canonicalization.
From empirical perspective: To further demonstrate how this discontinuity affects the accuracy of tensor property predictions, we conducted a set of comparative experiments in which we replaced the polar decomposition in GoeCTP with QR decomposition [22] for evaluation. In this QR decomposition case, the lattice matrix is decomposed as $\mathbf{L}=\mathbf{Q}\mathbf{R}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{R}$ is an upper triangular matrix that serves as the canonicalization norm. However, $\mathbf{R}$ does not satisfy Proposition A.3, and therefore does not constitute a continuous canonicalization approach.
To assess the sensitivity of each approach to small structural perturbations, we used the crystal $CdI_{2}$ (JARVIS_ID: JVASP-29631) as the model input, and then perturbed its lattice matrix (specifically, scaling down the first row of the lattice matrix) to observe the relative changes in the output (specifically the variations in the Fabenius norm of the output tensor). The detailed results are as follows:
Perturbation Ratio 0% 5% 10% 15% 20% 25% 30% 35% 40%
Output variation ratio (%): GoeCTP (QR) 0 3.38 6.63 11.29 21.83 39.37 67.81 124.79 191.28
Output variation ratio (%): GoeCTP (polar) 0 2.53 1.23 2.03 4.84 12.72 24.88 36.34 45.67
Table 24: Output variation ratio under different perturbation ratios on the dielectric dataset
The results clearly show that the output of GoeCTP using polar decomposition exhibits significantly smaller relative changes compared to the QR-based variant, particularly under increasing levels of perturbation. This highlights the robustness of the continuous canonicalization afforded by polar decomposition, demonstrating its superiority in maintaining prediction stability and structural consistency under small input variations.
A.15 How to design other canonical forms
In fact, the canonical form obtained via polar decomposition provides a foundation from which infinitely many alternative canonical forms can be derived. Specifically, given a crystal with lattice matrix $\mathbf{L}$, polar decomposition yields $\mathbf{L}=\mathbf{Q}\mathbf{H}$. While $\mathbf{H}$ serves as one valid canonical representation, additional canonical forms can be constructed by applying orthogonal transformations to $\mathbf{H}$. For instance, given a fixed orthogonal matrix $\mathbf{Q}_{0}$ (which may be arbitrarily chosen) or a neural network $f_{Q}(\mathbf{H})=\mathbf{Q}_{\theta}$ that predicts an orthogonal matrix $\mathbf{Q}_{\theta}$, one can construct new canonical forms such as $\mathbf{Q}_{0}\mathbf{H}$ or $\mathbf{Q}_{\theta}\mathbf{H}$. This is because, for any given crystal with an arbitrary orientation, a fixed canonical form can be obtained according to a predefined computation rule, i.e. $\mathbf{L}=\mathbf{Q}\mathbf{H}\quad\to\quad\mathbf{L}=\mathbf{Q}\mathbf{Q}_{0}^{\top}\mathbf{Q}_{0}\mathbf{H}\quad\text{or}\quad\mathbf{L}=\mathbf{Q}\mathbf{Q}_{\theta}^{\top}\mathbf{Q}_{\theta}\mathbf{H}$. This formulation highlights the inherent flexibility in defining canonical forms.
A.16 Impact of different canonical form on tensor property prediction
As discussed in Appendix A.15, there exist infinitely many possible canonical forms. To investigate the impact of different canonical forms on tensor property prediction in crystalline materials, we first present the results comparing the performance of GoeCTP with QR decomposition [22] and GoeCTP without canonicalization. The corresponding results are summarized in the table below:
GoeCTP (QR) GoeCTP (polar) GoeCTP (w/o QR or polar)
Fnorm $\downarrow$ 3.20 3.23 3.60
EwT 25% $\uparrow$ 83.5% 83.2% 80.6%
EwT 10% $\uparrow$ 56.0% 56.8% 56.2%
EwT 5% $\uparrow$ 32.4% 35.5% 32.6%
Total Time (s) $\downarrow$ 639 616 613
Table 25: Predictive performance comparisons between GoeCTP (QR), GoeCTP (polar), and GoeCTP (w/o QR or polar) on the dielectric dataset.
According to the results, it is evident that incorporating a canonical form improves tensor prediction performance compared to omitting it. While different canonical forms lead to slight variations in performance, the overall impact remains relatively modest. One possible explanation is that canonical forms typically exhibit structured representations, such as the symmetric positive semi-definite matrix obtained via polar decomposition or the upper triangular matrix from QR decomposition, which may facilitate the model’s ability to capture relevant features.
Furthermore, the use of a canonical form aligns all atoms within a crystal structure to a unified global coordinate system, thereby enabling equivariant tensor prediction. However, recent studies [43, 23, 15] have demonstrated that employing adaptive local frames, i.e. constructed based on individual atomic environments, can further enhance the accuracy of property predictions. Motivated by these insights, we extend the GoeCTP framework by incorporating local frames [43] (also referred to as local canonicalization). It is important to emphasize that GoeCTP can be seamlessly integrated with any prediction network, including those based on local frames, while preserving its equivariant prediction capability. The corresponding results are presented below.
GoeCTP (local frame+polar) GoeCTP (polar)
Fnorm $\downarrow$ 3.11 3.23
EwT 25% $\uparrow$ 83.9% 83.2%
EwT 10% $\uparrow$ 58.4% 56.8%
EwT 5% $\uparrow$ 33.5% 35.5%
Total Time (s) $\downarrow$ 1979 616
Table 26: Predictive performance comparisons between GoeCTP (local frame+polar) and GoeCTP (polar) on the dielectric dataset.
The results indicate that incorporating local frames can enhance predictive performance. However, this improvement comes at the expense of increased computational cost.